OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -3/5 and point (0, 0). A. y – 0 = -(3/5)(x – 0) B. y + 0 = (3/5)(x – 0) C. y + 1 = -(3/5)(x – 0) D. y – 1 = -(3/5)(x – 0)

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (tester97):

A. y – 0 = -(3/5)(x – 0)